OpenStudy (anonymous):

What is the simplified form of the expression √1/144

OpenStudy (shamil98):

what is the square root of 144?

OpenStudy (shamil98):

what is the square root of 1?

OpenStudy (shamil98):

\[\huge \frac{ \sqrt{1} }{ \sqrt{144} }\]

OpenStudy (anonymous):

12 and 1

OpenStudy (shamil98):

so your answer is?

OpenStudy (anonymous):

1/12...?

OpenStudy (shamil98):

yes